OpenStudy (anonymous):

Prove that if p and p²+8 are prime then so is p³+4.

OpenStudy (mitu12):

Every prime p, except 2 and 3 is congruent to 1 or 5 by modulo 6: p ≡ ±1 (mod 6) implies p² ≡ 1 (mod 6), p² + 8 ≡ 9 ≡ 3 (mod 6), so p² + 8 is prime only for p = 3. But then p³ + 4 = 31 - also prime.

OpenStudy (ikram002p):

how u get to know that p =1 or -1 mod 6 ?

OpenStudy (anonymous):

bbecause it is always positive

ganeshie8 (ganeshie8):

use division algorithm @ikram002p ?

OpenStudy (ikram002p):

ohhh ic , that make sense

ganeshie8 (ganeshie8):

2 | 6k+2 3 | 6k+3 2 | 6k+4

OpenStudy (anonymous):

can't be 2, 3, 4, mod six can it?

OpenStudy (ikram002p):

still , not sure of the proof i mean u can proof its its of the form 6k+1 or 6k+5 but how its possible to make sure its prime ?

ganeshie8 (ganeshie8):

its a necessary condition, not a sufficient condition

OpenStudy (ikram002p):

wait ok ok i got it nw hehehe

OpenStudy (anonymous):

Since all primes besides 2 are odd, and since 6k+1, 6k+5 are odd, there will be instances where these two forms generate primes. We're not saying that 6k+1 and 6k+5 are *always* prime.

OpenStudy (ikram002p):

ok the idea is p^2+8 is only prime when p=3 so we only need to check for p =3 wew ! i freaked out to see proof for prime forfirst look xD

OpenStudy (xapproachesinfinity):

what is modulo thing? I haven't seen that before

OpenStudy (ikram002p):

well not mudulo , try to think that p^2+8 is only prime when p=3 how could u show that ?

OpenStudy (ikram002p):

know about division algorithm ?

OpenStudy (xapproachesinfinity):

I'm coming back for this one lol, I haven't put my hand on this for a long time see ya, please leave this open

OpenStudy (ikram002p):

use D.A for P=3k+q since we need prime check for p=3k+1 and p=3k+2 when p=3k+1 then (3k+1)^2+8=9k^2+6k+9=3(,,,,) not prime for when p=3k+2 then (3k+2)^2+8=9k^2+12k+12=3(,,,,) not prime so only possible values p=2 or 3 but when its 2 then p^2+8 is not prime then p=3 when p=3 then p^2+8 and p³+4 both primes

OpenStudy (ikram002p):

@xapproachesinfinity u closed the Qn already :P

ganeshie8 (ganeshie8):

nice :) wonder if there is a way to prove this with out using DA

OpenStudy (ikram002p):

i guess there is hehe , anything leads to composite xD

OpenStudy (ikram002p):

mitu used the modulo thingy mmm

OpenStudy (ikram002p):

also since we know P^2+8 divides 3 for all p>3 then that should lead to something ,,,mmm like showing that p^2+8 is a product of 3 consecutive numbers ( don't know how this is gonna work though xD)

ganeshie8 (ganeshie8):

yeah DA is good enough i guess :)

ganeshie8 (ganeshie8):

i think we can't model a prime number w/o using DA

OpenStudy (ikram002p):

well mmm

OpenStudy (kainui):

Can we look at any tautologies to help us prove this? A= p is prime B= p²+8 is prime C= p³+4 is prime So we're saying this right? \[\Large A \wedge B \implies C\] This is the same as saying any of the following three I believe?: \[\Large A \implies (B \implies C) \\ \Large \overline C \implies \overline{(A \wedge B) } \\ \Large \overline C \implies \overline A \vee \overline B \\ \Large \] I wish I could offer something more helpful than this but I have a feeling that I'm using these incorrectly somehow, since I rarely do proofs.

ganeshie8 (ganeshie8):

last two statemetns are easy to see as they rely on demorgan laws

OpenStudy (ikram002p):

well :D u set up steps to proof xD ( in general )

ganeshie8 (ganeshie8):

first statement requires some thought + paper work lol

ganeshie8 (ganeshie8):

i think kainui is suggesting to prove it by contrapositive @ikram002p

OpenStudy (ikram002p):

use truth table to proof first one

OpenStudy (ikram002p):

yep yep

OpenStudy (ikram002p):

its like what we did in the proof @ganeshie8 xD we proved p^2+8 is composite

OpenStudy (xapproachesinfinity):

I'm sure of one thing lol, I only understood the statement that Kainui wrote down

OpenStudy (ikram002p):

mmm what ur course ? that asked u to prove this ? or what the subject ?

OpenStudy (xapproachesinfinity):

i was not the one who asked the question, I was just following lol, my course is just precalc but i had studied all this stuff long ago haha, i'm trying to get things back

OpenStudy (ikram002p):

ic .. well try to review division algorithm :) u will got this easily

OpenStudy (xapproachesinfinity):

Yeah I should go and review those stuff to better understand this